# Concentration Inequalities for Suprema of Empirical Processes with Dependent Data via Generic Chaining with Applications to Statistical Learning

### v1

@@ 3.1 Single-layer Neural Network
$\mathbb{P}\left(\left|{1\over n}\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta}_{1})-g(\bm{W}^{*}_{i,j},\bm{\theta}_{2})\right|\geq\left({C^{\prime}_{\alpha}C_{\Theta}\sqrt{\varepsilon}\over\sqrt{n}}+{C^{\prime\prime}_{\alpha}C_{\Theta}\varepsilon^{1\over\alpha}\over n^{1/\alpha\vee 1}}\right)d(\bm{\theta}_{1},\bm{\theta}_{2})\right)\leq e\exp\left(-\varepsilon\right)~.$
@@
Let $\{X_{\bm{\theta}}\}_{\bm{\theta}\in\Theta}$ be a separable zero-mean stochastic process on a metric space $(\Theta,d_{\Theta})$ that satisfies for any $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ in $\Theta$, some $\alpha>0$ and any $\varepsilon\geq 0$
@@
To simplify exposition, here we assume that $\Theta$ is finite (Talagrand, 2005, Ch. 2). [^3] We remark that Proposition 4.3 does not rely on this assumption and allows $\Theta$ to be uncountable. We begin by constructing a sequence of subsets of $\Theta$ denoted by $\{\Theta_{k}\}_{k\geq 0}^{K}$ such that $\bm{\theta}_{0}\in\Theta_{0}$ and $\Theta=\Theta_{K}$. The sequence of subsets is carefully constructed and may be interpreted as a sequence of progressively finer approximations of $\Theta$, in the sense that any $\bm{\theta}$ can be more accurately approximated by an element in $\Theta_{k}$ as $k$ increases. Let $\pi_{k}(\bm{\theta})=\arg\min_{s\in\Theta_{k}}d(s,\bm{\theta})$ denote the closest element of the set $\Theta_{k}$ to $\bm{\theta}$. Then, by constructing a telescoping sum and applying the triangle inequality we get that
@@
{ $\left\{\text{ for all }k\in\{1,\ldots,K\},\text{ for any }\bm{\theta}_{1},\bm{\theta}_{2}\in\Theta_{k},~|X_{\bm{\theta}_{1}}-X_{\bm{\theta}_{2}}|\leq c_{k}(\varepsilon)d(\bm{\theta}_{1},\bm{\theta}_{2})~\right\}~,$
@@
$\sup_{\bm{\theta}\in\Theta}|X_{\bm{\theta}}-X_{\bm{\theta}_{0}}|\leq\sup_{\bm{\theta}\in\Theta}\sum_{k\geq 1}^{K}|X_{\pi_{k}(\bm{\theta})}-X_{\pi_{k-1}(\bm{\theta})}|\leq\sup_{\bm{\theta}\in\Theta}\sum_{k\geq 1}^{K}c_{k}(\varepsilon)d(\pi_{k}(\bm{\theta}),\pi_{k-1}(\bm{\theta}))~.$
@@
Condition A.1theasm.(i)(i), Proposition 4.2 and Proposition 4.3 imply that for any $j\in\{0,\ldots,M-1\}$ and any $\varepsilon_{2}\geq 2$ we have that
$\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|\sum_{i=0}^{n}g(\bm{W}^{*}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta}_{0})\right|\geq 8{C^{\prime}_{\alpha}C_{\Theta}\gamma_{2}(\Theta)\sqrt{(\varepsilon^{\prime}_{2})}\over\sqrt{n}}+4^{(\alpha+1)/\alpha}{C^{\prime\prime}_{\alpha}C_{\Theta}\gamma_{\alpha}(\Theta)(\varepsilon^{\prime}_{2})^{1\over\alpha}\over n^{1/\alpha\vee 1}}\right)$
@@
$\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|{1\over n}\sum_{i=0}^{n}g(\bm{W}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta})\right|\geq{\varepsilon^{\prime}_{3}}\right)\leq\mathbb{P}\left({1\over n}\sum_{i=0}^{n}\sup_{\bm{\theta}\in\Theta}\left|g(\bm{W}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta})\right|\geq{\varepsilon^{\prime}_{3}}\right)$
$\quad\leq{1\over\varepsilon^{\prime}_{3}}{1\over n}\sum_{i=0}^{n}\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|g(\bm{W}_{i,j},\bm{\theta})-g(\bm{W}^{*}_{i,j},\bm{\theta})\right|\right)$
@@ 5 Conclusion
This paper establishes a concentration inequality for the suprema of the empirical processes with dependent data. The concentration inequality is established by developing an argument based on generic chaining combined with a coupling strategy. We apply our result to study the properties of statistical learning procedures. Specifically, we derive non-asymptotic predictive performance guarantees for empirical risk minimization for nonlinear regression. We show that empirical risk minimization achieves the classical convergence rate that can be obtained in i.i.d. setting after replacing the sample size what we call in this work the effective sample size, a notion of sample size that reflects the loss of information due to the dependence with respect to the i.i.d. case. Our result encompasses a broad class of nonlinear regression models, including a single-layer neural network models, and offers theoretical guarantees for widely used statistical learning procedures in dependent data environments.
@@ A.1 Proofs for Section 3
$(Y_{t}-f_{\bm{\theta}_{1}}(\bm{X}_{t}))^{2}-(Y_{t}-f_{\bm{\theta}_{2}}(\bm{X}_{t}))^{2}=(2Y_{t}-f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t}))(f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t}))$
$\quad\leq 2(|Y_{t}|+\sup_{\bm{\theta}}|f(\bm{X}_{t},\bm{\theta})|)|f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t})|~.$
Then C.1thecond and basic properties of subGaussian random variables imply that $\|g(\bm{Z}_{t},\bm{\theta}_{1})-g(\bm{Z}_{t},\bm{\theta}_{2})\|_{\psi_{1}}\leq 4C^{2}_{1}d(\bm{\theta}_{1},\bm{\theta}_{2})$. For any $\bm{\theta}\in\Theta$ we have
$g(\bm{Z}_{t},\bm{\theta})=(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}\leq 2Y_{t}^{2}+2f_{\bm{\theta}}(\bm{X}_{t})^{2}\leq 2(Y_{t}^{2}+\sup_{\bm{\theta}}f_{\bm{\theta}}(\bm{X}_{t})^{2})~.$
Then C.1thecond and Proposition A.3 imply that $\|g(\bm{Z}_{t},\bm{\theta}_{1})-\mathbb{E}g(\bm{Z}_{t},\bm{\theta}_{1})\|_{\psi_{1}}\leq 16C_{1}^{2}$. Thus A.1theasm is satisfied for $C_{\Theta}=16C_{1}^{2}$.
@@
$\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|g(\bm{Z}_{t},\bm{\theta})-g(\bm{Z}_{t}^{*},\bm{\theta})\right|\right)$
$\quad=\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t})-Y_{t}^{*}+f_{\bm{\theta}}(\bm{X}_{t}^{*}))+(Y_{t}^{*}-f_{\bm{\theta}}(\bm{X}_{t}^{*}))^{2}-(Y_{t}^{*}-f_{\bm{\theta}}(\bm{X}^{*}))^{2}\right|\right)$
@@
$\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|g(\bm{Z}_{t},\bm{\theta})-g(\bm{Z}_{t}^{*},\bm{\theta})\right|\right)\leq 12C_{1}\left\||Y_{t}-Y_{t}^{*}|+d_{\mathcal{X}}(\bm{X}_{t},\bm{X}_{t}^{*})\right\|_{L_{2}}~.$
Next define $d_{\mathcal{Z}}(\bm{z}_{1},\bm{z}_{2})=12C_{1}(|y_{1}-y_{2}|+d_{\mathcal{X}}(\bm{x}_{1},\bm{x}_{2}))$ and note that $d_{\mathcal{Z}}$ is a distance, since it is a sum of distances rescaled by a positive constant. Moreover, it is straightforward to verify that $d_{\mathcal{Z}}$ is also separable and complete which implies that $(\mathcal{Z},d_{\mathcal{Z}})$ is Polish. Moreover, if we pick $\bm{z}=(0,\bm{x})$ where $\bm{x}$ is is defined in C.2thecond we have that for any $s>0$ it holds that
@@
$13{T\over n}\exp(-\varepsilon)=13{1\over n}\text{ and }\beta^{1/4}\left(\left\lfloor{T\over n+1}\right\rfloor\right)\exp(\varepsilon)\leq n^{\zeta/4-\zeta/(4\eta)+1/\eta}\leq n^{-1/2}~.$
@@
We verify that C.1thecond and C.2thecond hold for the single-layer neural network. The claim then follows from Corollary 3.1.
@@
where, $\sigma(\bm{X}_{t},\bm{w}_{1},\ldots,\bm{w}_{K})=(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{1})\ldots\sigma(\bm{X}_{t}^{\prime}\bm{w}_{K}))^{\prime}$ and $\bm{\psi}=(\psi_{1},\ldots,\psi_{K})^{\prime}$. Second, since $\sigma$ has bounded first sub-derivative, it follows that $\sigma$ is Lipschitz and we shall the denote its Lipschitz constant by $L$. Third, since $\Theta$ is a compact set we have that we can find a positive constant $C_{\Theta}$ such that $\sup_{k}|\psi_{k}|\leq C_{\Theta}$ and $\sup_{k}\|\bm{w}_{k}\|_{2}\leq C_{\Theta}$. Fourth, since $\sigma$ is Lipschitz with Lipschitz constant $L$, $\sigma(0)=0$ and $\bm{X}_{t}$ is subGaussian with $\|\bm{X}_{t}\|_{\psi_{2}}<\sigma_{X}$, it follows that for all $k=1,\ldots,K$ we have that $\|\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})\|_{\psi_{2}}\leq C^{\prime}LC_{\Theta}\sigma_{X}$ where $C^{\prime}$ is a positive constant. To see this, note that if $\bm{Y}_{t}$ is an independent copy of $\bm{X}_{t}$ we have that, for any positive constant $c$,
$\mathbb{E}\exp{(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\mathbb{E}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k}))^{2}\over c^{2}}=\mathbb{E}\exp{(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\mathbb{E}(\sigma(\bm{Y}_{t}^{\prime}\bm{w}_{k})\mid\bm{X}_{t}))^{2}\over c^{2}}$
$\quad\leq\mathbb{E}\left(\mathbb{E}\left(\left.\exp{(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\sigma(\bm{Y}_{t}^{\prime}\bm{w}_{k}))^{2}\over c^{2}}\right|\bm{X}_{t}\right)\right)\leq\mathbb{E}\exp{L^{2}(\bm{X}_{t}^{\prime}\bm{w}_{k}-\bm{Y}_{t}^{\prime}\bm{w}_{k})^{2}\over c^{2}}$
$\quad=\mathbb{E}\exp{L^{2}\|\bm{w}_{k}\|^{2}_{2}(\bm{X}_{t}^{\prime}\bm{v}-\bm{Y}_{t}^{\prime}\bm{v})^{2}\over c^{2}}\leq\mathbb{E}\exp{2L^{2}\|\bm{w}_{k}\|^{2}_{2}[(\bm{X}_{t}^{\prime}\bm{v})^{2}+(\bm{Y}_{t}^{\prime}\bm{v})^{2}]\over c^{2}}$
$\quad=\left(\mathbb{E}\exp{2L^{2}\|\bm{w}_{k}\|^{2}_{2}(\bm{X}_{t}^{\prime}\bm{v})^{2}\over c^{2}}\right)^{2}=\mathbb{E}\exp{4L^{2}\|\bm{w}_{k}\|^{2}_{2}(\bm{X}_{t}^{\prime}\bm{v})^{2}\over c^{2}}~,$
where $\bm{v}=\bm{w}_{k}/\|\bm{w}_{k}\|_{2}$, where remark that the first inequality follows from Jensen’s inequality. If we then set $c=2L\sigma_{X}\|\bm{w}_{k}\|_{2}$ we have that the expectation in the last expression is at most 2, implying that $\|\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\mathbb{E}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})\|_{\psi_{2}}=2L\sigma_{X}\|\bm{w}_{k}\|_{2}\leq 2LC_{\Theta}\sigma_{X}$. Moreover, since $\sigma(0)=0$ it holds that
@@
$|f_{\bm{\theta}_{1}}(\bm{X})-f_{\bm{\theta}_{2}}(\bm{X})|=\left|\sum_{k=1}^{K}\psi_{1k}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{1k})-\psi_{2k}\sigma(\bm{X}^{\prime}\bm{w}_{2k})\right|$
@@
$\left\|\bm{X}_{t}^{\prime}{(\bm{w}_{1k}-\bm{w}_{2k})\over\|\bm{w}_{1k}-\bm{w}_{2k}\|_{2}}\right\|_{\psi_{2}}\leq\left\|\bm{X}_{t}\right\|_{\psi_{2}}<\sigma_{X}~.$
@@
$\|f_{\bm{\theta}_{1}}(\bm{X})-f_{\bm{\theta}_{2}}(\bm{X})\|_{\psi_{2}}\leq LC_{\Theta}\sigma_{X}\sum_{k=1}^{K}\|\bm{w}_{1k}-\bm{w}_{2k}\|_{2}+\sqrt{K}C^{\prime}LC_{\Theta}\sigma_{X}\|\bm{\psi}_{1}-\bm{\psi}_{2}\|_{2}$
@@
$\left\|\sup_{k}\sum_{k=1}^{K}\psi_{k}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})\right\|_{\psi_{2}}\leq\|\bm{\psi}\|_{2}\left\|\sigma(\bm{X}_{t},\bm{w}_{1},\ldots,\bm{w}_{k})^{\prime}{\bm{\psi}\over\|\bm{\psi}\|_{2}}\right\|_{\psi_{2}}\leq\sqrt{K}C^{\prime}LC_{\Theta}^{2}\sigma_{X}~.$
Additionally, by definition, $\|Y_{t}\|_{\psi_{2}}\leq\sigma_{Y}$. These, results confirm the second part of C.1thecond.
@@
$(i)$ The fact that $(\mathcal{X},d_{\mathcal{X}})$ is Polish follows from the properties of the maximum norm.
@@
The claim of the corrollary follows after noting that (Vershynin, 2026, Corollary 4.2.11) implies
@@
$\gamma_{2}(\Theta)\leq 2(\log(2))^{-1/2}\left(1-1/2^{1/2}\right)\sqrt{p}\Delta(\Theta)\leq 2(\log(2))^{-1/2}\left(1-1/2^{1/2}\right)\sqrt{2K}\sqrt{d}\Delta(\Theta)~.$

### v2

@@ 3.1 Single-layer Neural Network
$\mathbb{P}\left(\left|{1\over n}\sum_{i=0}^{n}g_{\bm{\theta}_{1}}(\bm{W}^{*}_{i,j})-g_{\bm{\theta}_{2}}(\bm{W}^{*}_{i,j})\right|\geq\left({C^{\prime}_{\alpha}C_{\Theta}\sqrt{\varepsilon}\over\sqrt{n}}+{C^{\prime\prime}_{\alpha}C_{\Theta}\varepsilon^{1\over\alpha}\over n^{1/\alpha\vee 1}}\right)d_{\Theta}(\bm{\theta}_{1},\bm{\theta}_{2})\right)\leq e\exp\left(-\varepsilon\right)~.$
@@
Let $\{X_{\bm{\theta}}\}_{\bm{\theta}\in\Theta}$ be a separable zero-mean stochastic process on a metric space $(\Theta,d_{\Theta})$ that satisfies for any $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ in $\Theta$, some $a,b,\alpha>0$ and any $\varepsilon\geq 0$
@@
To simplify exposition, here we assume that $\Theta$ is finite (Talagrand, 2005, Ch. 2). [^3] We remark that Proposition 4.3 does not rely on this assumption and allows $\Theta$ to be uncountable. We begin by constructing a sequence of subsets of $\Theta$ denoted by $\{\Theta_{k}\}_{k\geq 0}^{K}$ such that $\bm{\theta}_{0}\in\Theta_{0}$ and $\Theta=\Theta_{K}$. The sequence of subsets is carefully constructed and may be interpreted as a sequence of progressively finer approximations of $\Theta$, in the sense that any $\bm{\theta}$ can be more accurately approximated by an element in $\Theta_{k}$ as $k$ increases. Let $\pi_{k}(\bm{\theta})=\arg\min_{s\in\Theta_{k}}d_{\Theta}(s,\bm{\theta})$ denote the closest element of the set $\Theta_{k}$ to $\bm{\theta}$. Then, by constructing a telescoping sum and applying the triangle inequality we get that
@@
{ $\left\{\text{ for all }k\in\{1,\ldots,K\},\text{ for any }\bm{\theta}_{1},\bm{\theta}_{2}\in\Theta_{k},~|X_{\bm{\theta}_{1}}-X_{\bm{\theta}_{2}}|\leq c_{k}(\varepsilon)d_{\Theta}(\bm{\theta}_{1},\bm{\theta}_{2})~\right\}~,$
@@
$\sup_{\bm{\theta}\in\Theta}|X_{\bm{\theta}}-X_{\bm{\theta}_{0}}|\leq\sup_{\bm{\theta}\in\Theta}\sum_{k\geq 1}^{K}|X_{\pi_{k}(\bm{\theta})}-X_{\pi_{k-1}(\bm{\theta})}|\leq\sup_{\bm{\theta}\in\Theta}\sum_{k\geq 1}^{K}c_{k}(\varepsilon)d_{\Theta}(\pi_{k}(\bm{\theta}),\pi_{k-1}(\bm{\theta}))~.$
@@
Condition A.1theasm.(i)(i), Proposition 4.2 and Proposition 4.3 imply that for any $j\in\{0,\ldots,M-1\}$ and any $\varepsilon^{\prime}_{2}\geq 2$ we have that
$\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}^{*}_{i,j})-g_{\bm{\theta}_{0}}(\bm{W}^{*}_{i,j})\right|\geq 8{C^{\prime}_{\alpha}C_{\Theta}\gamma_{2}(\Theta)\sqrt{(\varepsilon^{\prime}_{2})}\over\sqrt{n}}+4^{(\alpha+1)/\alpha}{C^{\prime\prime}_{\alpha}C_{\Theta}\gamma_{\alpha}(\Theta)(\varepsilon^{\prime}_{2})^{1\over\alpha}\over n^{1/\alpha\vee 1}}\right)$
@@
$\mathbb{P}\left(\sup_{\bm{\theta}\in\Theta}\left|{1\over n}\sum_{i=0}^{n}g_{\bm{\theta}}(\bm{W}_{i,j})-g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{3}}\right)\leq\mathbb{P}\left({1\over n}\sum_{i=0}^{n}\sup_{\bm{\theta}\in\Theta}\left|g_{\bm{\theta}}(\bm{W}_{i,j})-g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|\geq{\varepsilon^{\prime}_{3}}\right)$
$\quad\leq{1\over\varepsilon^{\prime}_{3}}{1\over n}\sum_{i=0}^{n}\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|g_{\bm{\theta}}(\bm{W}_{i,j})-g_{\bm{\theta}}(\bm{W}^{*}_{i,j})\right|\right)$
@@ 5 Conclusion
This paper establishes a concentration inequality for the suprema of the empirical processes with dependent data. The concentration inequality is established by developing an argument based on generic chaining combined with a coupling strategy. We apply our result to study the properties of statistical learning procedures. Specifically, we derive non-asymptotic predictive performance guarantees for empirical risk minimization for nonlinear regression. We show that empirical risk minimization achieves the classical convergence rate that can be obtained in i.i.d. setting after replacing the sample size with what we call in this work the effective sample size, a notion of sample size that reflects the loss of information due to the dependence with respect to the i.i.d. case. Our result encompasses a broad class of nonlinear regression models, including a single-layer neural network models, and offers theoretical guarantees for widely used statistical learning procedures in dependent data environments.
@@ A.1 Proofs for Section 3
$(Y_{t}-f_{\bm{\theta}_{1}}(\bm{X}_{t}))^{2}-(Y_{t}-f_{\bm{\theta}_{2}}(\bm{X}_{t}))^{2}=(-2Y_{t}+f_{\bm{\theta}_{1}}(\bm{X}_{t})+f_{\bm{\theta}_{2}}(\bm{X}_{t}))(f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t}))$
$\quad\leq 2(|Y_{t}|+\sup_{\bm{\theta}}|f_{\bm{\theta}}(\bm{X}_{t})|)|f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t})|~.$
Then C.1thecond and basic properties of subGaussian random variables imply that $\|g_{\bm{\theta}_{1}}(\bm{Z}_{t})-g_{\bm{\theta}_{2}}(\bm{Z}_{t})\|_{\psi_{1}}\leq 4C^{2}_{1}d_{\Theta}(\bm{\theta}_{1},\bm{\theta}_{2})$. For any $\bm{\theta}\in\Theta$ we have
$g_{\bm{\theta}}(\bm{Z}_{t})=(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))^{2}\leq 2Y_{t}^{2}+2f_{\bm{\theta}}(\bm{X}_{t})^{2}\leq 2(Y_{t}^{2}+\sup_{\bm{\theta}}f_{\bm{\theta}}(\bm{X}_{t})^{2})~.$
Then C.1thecond and Proposition A.3 imply that $\|g_{\bm{\theta}}(\bm{Z}_{t})-\mathbb{E}g_{\bm{\theta}}(\bm{Z}_{t})\|_{\psi_{1}}\leq 16C_{1}^{2}$. Thus A.1theasm is satisfied for $C_{\Theta}=16C_{1}^{2}$.
@@
$\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|g_{\bm{\theta}}(\bm{Z}_{t})-g_{\bm{\theta}}(\bm{Z}_{t}^{*})\right|\right)$
$\quad=\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t})-Y_{t}^{*}+f_{\bm{\theta}}(\bm{X}_{t}^{*}))^{2}+2(Y_{t}^{*}-f_{\bm{\theta}}(\bm{X}_{t}^{*}))[(Y_{t}-f_{\bm{\theta}}(\bm{X}_{t}))-(Y_{t}^{*}-f_{\bm{\theta}}(\bm{X}_{t}^{*}))]\right|\right)$
@@
$\mathbb{E}\left(\sup_{\bm{\theta}\in\Theta}\left|g_{\bm{\theta}}(\bm{Z}_{t})-g_{\bm{\theta}}(\bm{Z}_{t}^{*})\right|\right)\leq 12C_{1}\left\||Y_{t}-Y_{t}^{*}|+d_{\mathcal{X}}(\bm{X}_{t},\bm{X}_{t}^{*})\right\|_{L_{2}}~.$
Next define $d_{\mathcal{Z}}(\bm{z}_{1},\bm{z}_{2})=12C_{1}(|y_{1}-y_{2}|+d_{\mathcal{X}}(\bm{x}_{1},\bm{x}_{2}))$ and note that $d_{\mathcal{Z}}$ is a distance, since it is a sum of distances rescaled by a positive constant. Moreover, it is straightforward to verify that the space $\mathcal{Z}$ equipped with the distance $d_{\mathcal{Z}}$ is also separable and complete which implies that $(\mathcal{Z},d_{\mathcal{Z}})$ is Polish. Since $(\mathcal{X},d_{\mathcal{X}})$ is Polish by assumption, it suffices to prove that $(\mathcal{Y},|\cdot|)$ is Polish as well; in that case, the product space $(\mathcal{Z},d_{\mathcal{Z}})$ is automatically Polish. Recall that $\mathcal{Y}\subset\mathbb{R}$ is closed. Equipped with the metric induced by the Euclidean distance on $\mathbb{R}$, any subset of $\mathbb{R}$ is separable, and any closed subset is complete. It follows that $(\mathcal{Y},|\cdot|)$ is a Polish space. If we pick $\bm{z}=(0,\bm{x})$ where $\bm{x}$ is is defined in C.2thecond we have that for any $s>0$ it holds that
@@
$13{T\over n}\exp(-\varepsilon)=13{1\over n}\text{ and }\beta^{1/4}\left(\left\lfloor{T\over n+1}\right\rfloor\right)\exp(\varepsilon)\leq n^{\zeta/4-\zeta/(4\eta)+1/\eta}=n^{-1/2}~,$
having used that $\eta=\frac{\zeta-4}{\eta+2}$. We remark that the conditions of the Theorem require $\zeta>4$ so that $\eta>0$ and consequently $\varepsilon>0$. ∎
@@
We verify that C.1thecond and C.2thecond hold for the single-layer neural network. The claim then follows from Proposition 3.1.
@@
where, $\sigma(\bm{X}_{t},\bm{w}_{1},\ldots,\bm{w}_{K})=(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{1}),\ldots,\sigma(\bm{X}_{t}^{\prime}\bm{w}_{K}))^{\prime}$ and $\bm{\psi}=(\psi_{1},\ldots,\psi_{K})^{\prime}$. Second, since $\sigma$ has bounded first sub-derivative, it follows that $\sigma$ is Lipschitz and we shall denote its Lipschitz constant by $L$. Third, since $\Theta$ is a compact set we have that we can find a positive constant $C_{\Theta}$ such that $\sup_{k}|\psi_{k}|\leq C_{\Theta}$ and $\sup_{k}\|\bm{w}_{k}\|_{2}\leq C_{\Theta}$. Fourth, since $\sigma$ is Lipschitz with Lipschitz constant $L$, $\sigma(0)=0$ and $\bm{X}_{t}$ is sub-Gaussian with $\|\bm{X}_{t}\|_{\psi_{2}}=\sigma_{X}$, it follows that for all $k=1,\ldots,K$ we have that $\|\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})\|_{\psi_{2}}\leq C^{\prime}LC_{\Theta}\sigma_{X}$ where $C^{\prime}$ is a positive constant. To see this, note that if $\widetilde{\bm{X}}_{t}$ is an independent copy of $\bm{X}_{t}$ we have that, for any positive constant $c$,
$\mathbb{E}\exp{(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\mathbb{E}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k}))^{2}\over c^{2}}=\mathbb{E}\exp{(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\mathbb{E}(\sigma(\widetilde{\bm{X}}_{t}^{\prime}\bm{w}_{k})\mid\bm{X}_{t}))^{2}\over c^{2}}$
$\quad\leq\mathbb{E}\left(\mathbb{E}\left(\left.\exp{(\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\sigma(\widetilde{\bm{X}}_{t}^{\prime}\bm{w}_{k}))^{2}\over c^{2}}\right|\bm{X}_{t}\right)\right)\leq\mathbb{E}\exp{L^{2}(\bm{X}_{t}^{\prime}\bm{w}_{k}-\widetilde{\bm{X}}_{t}^{\prime}\bm{w}_{k})^{2}\over c^{2}}$
$\quad=\mathbb{E}\exp{L^{2}\|\bm{w}_{k}\|^{2}_{2}(\bm{X}_{t}^{\prime}\bm{v}-\widetilde{\bm{X}}_{t}^{\prime}\bm{v})^{2}\over c^{2}}\leq\mathbb{E}\exp{2L^{2}\|\bm{w}_{k}\|^{2}_{2}[(\bm{X}_{t}^{\prime}\bm{v})^{2}+(\widetilde{\bm{X}}_{t}^{\prime}\bm{v})^{2}]\over c^{2}}$
$\quad=\left(\mathbb{E}\exp{2L^{2}\|\bm{w}_{k}\|^{2}_{2}(\bm{X}_{t}^{\prime}\bm{v})^{2}\over c^{2}}\right)^{2}\leq\mathbb{E}\exp{4L^{2}\|\bm{w}_{k}\|^{2}_{2}(\bm{X}_{t}^{\prime}\bm{v})^{2}\over c^{2}}~,$
where $\bm{v}=\bm{w}_{k}/\|\bm{w}_{k}\|_{2}$ and we remark that the first inequality follows from Jensen’s inequality. If we then set $c=2L\sigma_{X}\|\bm{w}_{k}\|_{2}$ we have that the expectation in the last expression is at most 2, implying that $\|\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})-\mathbb{E}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})\|_{\psi_{2}}=2L\sigma_{X}\|\bm{w}_{k}\|_{2}\leq 2LC_{\Theta}\sigma_{X}$. Moreover, since $\sigma(0)=0$ it holds that
@@
$|f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t})|=\left|\sum_{k=1}^{K}\psi_{1k}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{1k})-\psi_{2k}\sigma(\bm{X}^{\prime}\bm{w}_{2k})\right|$
@@
$\left\|\bm{X}_{t}^{\prime}{(\bm{w}_{1k}-\bm{w}_{2k})\over\|\bm{w}_{1k}-\bm{w}_{2k}\|_{2}}\right\|_{\psi_{2}}\leq\left\|\bm{X}_{t}\right\|_{\psi_{2}}=\sigma_{X}~.$
@@
$\|f_{\bm{\theta}_{1}}(\bm{X}_{t})-f_{\bm{\theta}_{2}}(\bm{X}_{t})\|_{\psi_{2}}\leq LC_{\Theta}\sigma_{X}\sum_{k=1}^{K}\|\bm{w}_{1k}-\bm{w}_{2k}\|_{2}+\sqrt{K}C^{\prime}LC_{\Theta}\sigma_{X}\|\bm{\psi}_{1}-\bm{\psi}_{2}\|_{2}$
@@
$\left\|\sum_{k=1}^{K}\psi_{k}\sigma(\bm{X}_{t}^{\prime}\bm{w}_{k})\right\|_{\psi_{2}}\leq\|\bm{\psi}\|_{2}\left\|\sigma(\bm{X}_{t},\bm{w}_{1},\ldots,\bm{w}_{K})^{\prime}{\bm{\psi}\over\|\bm{\psi}\|_{2}}\right\|_{\psi_{2}}\leq KC^{\prime}LC_{\Theta}^{2}\sigma_{X}~.$
Additionally, by definition, $\|Y_{t}\|_{\psi_{2}}=\sigma_{Y}$. These results confirm the second part of C.1thecond.
@@
$(i)$ The fact that $(\mathcal{X},d_{\mathcal{X}})$ is Polish follows from the fact that $\mathbb{R}^{d}$ equipped with the sup-norm is Polish, and any closed subset of a Polish space is Polish as well.
@@
The claim of the corollary follows after noting that (Vershynin, 2026, Corollary 4.2.11) implies
@@
$\gamma_{2}(\Theta)\leq 2(\log(2))^{-1/2}\left(1-1/2^{1/2}\right)\sqrt{p}\Delta(\Theta)\leq 2(\log(2))^{-1/2}\left(1-1/2^{1/2}\right)\sqrt{2K}\sqrt{d}\Delta(\Theta)~,$
where we recall that $\Delta(\Theta)$ is the diameter of the set $\Theta$.